# Gradient Field-Based Dynamic Window Approach for Collision Avoidance in Complex Environments

### v1

@@ I INTRODUCTION
Escaping from a non-convex obstacle requires more than just instantaneous distance information; it is essential to determine whether the mobile robot is approaching the obstacle. A straightforward solution is to leverage the gradient of the distance to the obstacle, allowing the robot to take actions that guide it along the gradient direction, thereby avoiding potential collisions. The Euclidean signed distance field [16, 17] is a widely used method to model the environment, as it provides both distance and gradient information relative to the nearest occupied cell. However, to capture the overall shape of the obstacle, a robot needs to consider the distance and gradient information from all directions, which is crucial for non-convex obstacle avoidance. The Gaussian Process Distance Field (GPDF) [18, 19] offers a probabilistic approach to modeling the environment, providing distance and gradient information from all detected obstacles. This enables a more comprehensive representation of non-convex structures, allowing for more effective motion planning and obstacle avoidance in complex environments.
@@ III-A Predictive Obstacle Avoidance based on Gradient Field
where $\Delta\theta=\theta^{(n)}-\theta^{(n)}_{d}$ is the difference between the heading angle $\theta^{(n)}$ of the robot and the heading direction $\theta^{(n)}_{d}$ of the gradient at state $\bm{x}^{(n)}$, $\Delta\theta_{\text{thre}}$ is the threshold to include the collision-gradient term, and $\beta$ is a tuning parameter to control changing rate of the cost term. The cost term $J_{\text{col-grad}}$ is designed to be small when the robot is moving away from the obstacle and large when the robot is moving toward the obstacle. Although the gradient-based term $J_{\text{col-grad}}$ can capture the potential collision risks, it may lead to overly conservative behavior, causing the robot to avoid obstacles even when not moving toward them. The choice of $\Delta\theta_{\text{thre}}$ plays a crucial role in balancing the trade-off between collision avoidance and regular path following. Consequently, it should be satisfies that $\Delta\theta_{\text{thre}}\in(\pi/2,\pi]$. Fig. 2 visualizes this cost term for a U-shaped non-convex obstacle. The distance field alone creates a local minimum near the dead end, while the gradient-based cost successfully identifies the directions leading to the dead end, preventing the robot from becoming trapped.
@@ IV Implementation
where $d$ is the distance between two points and $L=0.2$ is a positive parameter controlling the interpolation degree. The boundary points of obstacles are sampled from the obstacle edges in the resolution of 0.1 meters. For the DWA, the horizon $N$ is 20, the threshold to include the collision-gradient term is set as $\Delta\theta_{\text{thre}}=2\pi/3$, and the tuning parameter is set as $\beta=2$. In addition, to avoid over-conservative behavior, the collision-avoidance cost term is inactive unless any point on the candidate is within a one-meter range of any obstacle. To generate the candidate control set, the speed resolution is 0.3 meters per second and the angular velocity resolution is 0.08 radians per second (about 4.6 degrees per second). The maximal total number of candidate trajectories is 84.
@@ V Evaluation
In the evaluation [^1] The demo video is available at https://youtu.be/Et4C5oGHlF0, the proposed GF-DWA is compared with the regular DWA, optimization-based MPC, and learning-based Deep Deterministic Policy Gradient (DDPG) algorithm [20] in different scenarios. The regular DWA shares the same objective function as GF-DWA except for the collision-avoidance term, which is only based on the minimal distance to any obstacle in regular DWA. The implementation of MPC and DDPG is according to [13]. For the single-robot cases, the success rates of selected methods are compared. A successful run means the robot can reach the final goal without any collisions. For the multi-robot cases, besides the success rate, the number of time steps needed to finish the case is also compared. Note that the proposed GF-DWA is sampling-based, which does not require solving an optimization problem at each time step or pre-training any models.
@@ V-A Static Obstacle Avoidance
Figure 6: Evaluation of the multi-agent cases. Each bar shows the average time step to reach the goals. Black bars indicate the standard deviations of instances in the corresponding category. Black dots are individual trials. The timeout limit is 200 steps, which means MPC has failures in the first case.
V-B Multi-robot collision avoidance
@@
Unlike the single-robot static obstacle avoidance tests, all compared methods can handle the multi-robot tests but with varying levels of efficiency and stability. The detailed finishing time step of each trial is reported in Fig. 6. While DWA and GF-DWA generate deterministic control actions, MPC exhibits instability since it requires numerical solvers in practice. In multi-robot test 1, DWA completes the task faster than GF-DWA because it doesn’t take the predicted states of other robots into account and lacks proactive evasive behavior. As a result, the intersection distance between different robots is very small, posing a significant collision risk in practical deployment. On the contrary, GF-DWA produces predictive collision avoidance behavior as shown in Fig. 4, ensuring a safer and more stable navigation strategy. Overall, GF-DWA outperforms other methods by achieving a better balance between stability, safety, and efficiency, making it a more robust solution for multi-robot collision avoidance in complex environments.

### v2

@@ I INTRODUCTION
Escaping from a non-convex obstacle requires more than just instantaneous distance information; it is essential to determine whether the mobile robot is approaching the obstacle. A straightforward solution is to leverage the gradient of the distance to the obstacle, allowing the robot to take actions that guide it along the gradient direction, thereby avoiding potential collisions. The Euclidean Signed Distance Field (ESDF) [16, 17] is a widely used method to model the environment, as it provides both distance and gradient information relative to the nearest occupied cell. However, to capture the overall shape of the obstacle, a robot needs to consider the distance and gradient information from all directions, which is crucial for non-convex obstacle avoidance. The Gaussian Process Distance Field (GPDF) [18, 19] offers a probabilistic approach to modeling the environment, providing distance and gradient information from all detected obstacles. This enables a more comprehensive representation of non-convex structures, allowing for more effective motion planning and obstacle avoidance in complex environments.
@@ III-A Predictive Obstacle Avoidance based on Gradient Field
where $\Delta\theta=\theta^{(n)}-\theta^{(n)}_{d}$ is the difference between the heading angle $\theta^{(n)}$ of the robot and the normalized heading direction $\theta^{(n)}_{d}$ of the gradient at state $\bm{x}^{(n)}$, $\Delta\theta_{\text{thre}}$ is the threshold to include the collision-gradient term, and $\beta$ is a tuning parameter to control changing rate of the cost term. The cost term $J_{\text{col-grad}}$ is designed to provide a smooth, continuous cost increase as heading misalignment grows. The exponential form imposes stronger penalties when the robot moves toward obstacles in its forward direction. Although the gradient-based term $J_{\text{col-grad}}$ can capture the potential collision risks, it may lead to overly conservative behavior, causing the robot to avoid obstacles even when not moving toward them. The choice of $\Delta\theta_{\text{thre}}$ plays a crucial role in balancing the trade-off between collision avoidance and regular path following. Consequently, it should be satisfies that $\Delta\theta_{\text{thre}}\in(\pi/2,\pi]$. Fig. 2 visualizes this cost term for a U-shaped non-convex obstacle. The distance field alone creates a local minimum near the dead end, while the gradient-based cost successfully identifies the directions leading to the dead end, preventing the robot from becoming trapped.
@@ IV Implementation
where $d$ is the distance between two points and $L=0.2$ is a positive parameter controlling the interpolation degree. The boundary points of obstacles are sampled from the obstacle edges at a resolution of 0.1 meters. For the DWA, the horizon $N$ is 20, the threshold to include the collision-gradient term is set as $\Delta\theta_{\text{thre}}=2\pi/3$, and the tuning parameter is set as $\beta=2$. The choice of $\Delta\theta_{\text{thre}}$ creates a strong repulsive force when heading toward obstacles, and minimizes impact when moving parallel or away from obstacles. In addition, to avoid over-conservative behavior, the collision-avoidance cost term is inactive unless any point on the candidate is within a one-meter range of any obstacle. To generate the candidate control set, the speed resolution is 0.3 meters per second and the angular velocity resolution is 0.08 radians per second (about 4.6 degrees per second). The maximal total number of candidate trajectories is 84. For the GPDF, after ahead-of-time compilation, on an Intel i7-9750H CPU, the inference time is around 0.008 seconds in every time step, which achieves highly real-time performance. Note that this can be further improved using GPUs.
@@ V Evaluation
In the evaluation [^1] The demo video is available at https://youtu.be/Et4C5oGHlF0, the proposed GF-DWA is compared with the regular DWA, optimization-based MPC, and learning-based Deep Deterministic Policy Gradient (DDPG) algorithm [20] in different scenarios. The regular DWA shares the same objective function as GF-DWA except for the collision-avoidance term, which is only based on the minimal distance to any obstacles. The implementation of MPC and DDPG is according to [13]. For the single-robot cases, the success rates of selected methods are compared. A successful run means the robot can reach the final goal without any collisions. For the multi-robot cases, besides the success rate, the number of time steps needed to finish the case is also compared. Note that the proposed GF-DWA is sampling-based, with no need to solve an optimization problem at each time step or pre-train any models.
@@ V-A Static Obstacle Avoidance
Figure 6: Evaluation of the multi-agent cases. Each bar shows the average time step to reach the goals. Black bars indicate the standard deviations of instances in the category. Black dots are individual trials. The timeout limit is 200 steps, which means MPC has failures in the first case.
V-B Gradient information
To further investigate the effectiveness of the gradient information from GPDFs, additional tests were conducted in S3 with a U-shape obstacle. Variants of DWA were tested using distance-only data from GPDF, gradient data from ESDF, and gradient data from GPDF. Results show that only the GPDF-based gradient enables successful navigation, as ESDF gradients reflect only the nearest obstacle point and fail to capture the global structure of non-convex obstacles.
V-C Multi-robot collision avoidance
@@
Unlike the single-robot static obstacle avoidance tests, all compared methods can handle the multi-robot tests but with varying levels of efficiency and stability. The detailed finishing time step of each trial is reported in Fig. 6. While DWA and GF-DWA generate deterministic control actions, MPC exhibits instability since it requires numerical solvers in practice. In multi-robot test 1, for the regular DWA, the intersection distance between different robots is very small, posing a significant collision risk in practical deployment. On the contrary, GF-DWA produces predictive collision avoidance behavior as shown in Fig. 4, ensuring a safer and more stable navigation strategy. Overall, GF-DWA outperforms other methods by achieving a better balance between stability, safety, and efficiency, making it a more robust solution for multi-robot collision avoidance in complex environments.